# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v5

@@ Abstract
The rapid advancements in Pattern Analysis and Machine Intelligence (PAMI) have led to an overwhelming expansion of scientific knowledge, spawning numerous literature reviews aimed at collecting and synthesizing fragmented information. This paper presents a thorough analysis of these literature reviews within the PAMI field, and tries to address three core research questions: (1) What are the prevalent structural and statistical characteristics of PAMI literature reviews? (2) What strategies can researchers employ to efficiently navigate the growing corpus of reviews? (3) What are the advantages and limitations of AI-generated reviews compared to human-authored ones? To address the first research question, we begin with a narrative overview to highlight common preferences in composing PAMI reviews, followed by a statistical analysis to quantitatively uncover patterns in these preferences. Our findings reveal several key insights. First, fewer than 20% of PAMI reviews currently comply with PRISMA standards, although this proportion is gradually increasing. Second, there is a moderate positive correlation between the quality of references and the scholarly impact of reviews, emphasizing the importance of reference selection. To further assist researchers in efficiently managing the rapidly growing number of literature reviews, we introduce four novel, real-time, article-level bibliometric indicators that facilitate the screening of numerous reviews. Finally, our comparative analysis reveals that AI-generated reviews currently fall short of human-authored ones in accurately evaluating the academic significance of newly published articles and integrating rich visual elements, which limits their practical utility. Overall, this study provides a deeper understanding of PAMI literature reviews by uncovering key trends, evaluating current practices, and highlighting areas for future improvement.
Index Terms:
literature review, PAMI, umbrella review, AI-generated review, bibliometrics
@@ 1 Introduction
THE entropy of almost all natural and artificial systems in the universe exhibits a continuous increase. This principle extends to humanity’s knowledge system, which similarly experiences a steady rise in entropy. Such growing disorder within the knowledge system can lead to redundant efforts, hindering the efficient creation of new knowledge. Analogous to the role of gravity in shaping the early universe filled with diverse particles, the literature review serves as a critical mechanism for organizing dispersed knowledge within the academic sphere. Essentially, a literature review is a scholarly composition that not only demonstrates an understanding of the academic literature on a given topic but also situates this knowledge within a broader context. It compiles the most relevant and significant publications related to a specific research area, thereby offering a comprehensive overview of that field.
@@
Figure 1: Annual publication trends of literature review in the field of PAMI. A notable rising trend can be observed since 2015, reflecting an increasing scholarly focus and growing recognition of the importance of review articles in synthesizing the state of research in PAMI. The data is collected from the Google Scholar search engine.
@@ 1.1 Research Question and Contribution
This paper specifically focuses on reviews within the field of PAMI and tries to answer the following three core Research Questions (RQs):
RQ1: What are the prevalent structural and statistical features shared among literature reviews within the PAMI field? Such information helps researchers develop a more specific understanding of literature reviews, thereby advancing the field’s overall comprehension of the role and significance of these reviews.
RQ2: What strategy can researchers employ to navigate the expanding corpus of literature reviews in PAMI more efficiently? A suitable strategy facilitates readers in navigating a large number of similar reviews and identifying the reviews that best match their interests and objectives.
RQ3: What are the advantages and limitations of AI-generated reviews, and what are the key differences between AI-generated and human-authored literature reviews in PAMI? A comprehensive understanding of AI-generated reviews enables users to benefit from the convenience offered by automated review technologies while minimizing the risk of potential misuse.
To address these questions:
  • We conduct both narrative review and statistical analysis to investigate the writing styles, meta-data distributions, and features of literature reviews in the PAMI field. This information aims to help researchers develop a more in-depth understanding of the nature of literature reviews.
  • We propose four real-time, article-level, and field-normalized bibliometric indicators for the quantitative evaluation of literature reviews. It is believed that the integrated use of these indicators may help researchers make more informed choices when selecting review articles.
  • We evaluate several existing automated review generation systems and conducted a systematic assessment and bibliometric analysis of these AI-generated reviews. The findings highlighted notable differences between AI-generated reviews and those written by humans, particularly in the quality of references, and the richness of visual elements.
@@ 1.2 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 briefly introduces the criteria for literature screening, methods for the RiPAMI construction, and formulas for metric calculations. Section 3 provides a narrative analysis of the structure and content features of the selected reviews. Section 4 presents a statistical analysis of meta-data and content features of reviews in the RiPAMI. Section 5 further discusses how the proposed metrics can assist in efficiently selecting the proper reviews. The characteristics of human-authored versus AI-generated literature reviews are discussed in Sec. 6. In Sec. 7, we explore the challenges and future prospects of literature reviews. Finally, the paper concludes in Sec. 8.
@@ 2.1 Scope
This study examines all fourteen types of review articles as classified in Grant’s comprehensive typology [33], including narrative review, systematic review, state-of-the-art review, etc. While this research is technically a tertiary or “umbrella” study, we opted for the more commonly understood term “literature review” to improve clarity and acceptance within the PAMI field. In addition, although the quantity of reviews is considerably smaller compared to that of normal papers, it remains impractical to analyze reviews within every field. Therefore, this paper will focus only on reviews in the PAMI field.
2.2 Literature Selection Criteria for Classical Review
To identify the features necessary for statistical analysis, we conduct a classical narrative review to determine the key characteristics of literature reviews in the PAMI field. This review utilizes resources such as Google Scholar, Semantic Scholar, IEEE Xplore, and ScienceDirect, covering publications from major publishers including IEEE, Elsevier, Springer, and others. To ensure that the selected papers are both review articles and highly relevant to the PAMI domain, we adopt a straightforward and effective approach. First, the 106 search keywords are derived from the scopes of prominent journals and conferences in the PAMI field, including terms such as speech recognition, optical character recognition, and self-supervised learning. Second, a filtering criterion is applied: the paper’s title must include either “survey” or “review”, and the specified search keyword must appear in the title or abstract. Finally, to maintain the quality of the included reviews, pre-prints not published in journals or conference proceedings are excluded, ensuring that all selected articles are peer-reviewed. This approach helps to maximize the inclusion of high-quality and relevant literature for subsequent analysis.
@@ 2.3.1 Data Source
TABLE I: Comparisons between various sources: “Counts only” in the citations column means that the source only records the citation counts, while “Complete” signifies that the data source provides a complete list of citations.
@@
Table I compares various most commonly used data sources. However, none of these sources is perfect. For example, arXiv is a free distribution service and an open-access archive for millions of scholarly articles in various fields. However, the arXiv data source fails to provide the publication venue, citations, and references. Semantic Scholar seems promising, but it suffers a lower update rate than arXiv and a narrower search scope than Google Scholar. Google Scholar is an online search engine that indexes scholarly literature from a wide range of disciplines. Although Google Scholar uses powerful automated programs to retrieve files for inclusion in its search results, it still faces challenges in many areas. Beel [9, 10, 11] argues that Google Scholar places a high weight on citation counts in its ranking algorithm and has therefore been criticized for exacerbating the Matthew effect [83]. Moreover, the citation counts displayed on Google Scholar are subject to manipulation by complete nonsense articles indexed on Google Scholar (e.g. citations from AI-generated pre-print papers published on arXiv should have been ignored). Therefore, a promising engineering solution is to leverage the strengths of various approaches to overcome the weaknesses of each approach, as will be detailed in the next subsection.
@@ 2.3.2 Database construction
[…]
@@
Figure 2: A schematic overview of the database construction process. From initial keyword selection to the final SQL-based RiPAMI snapshot, three key steps are implemented to ensure the data in RiPAMI is clean, accurate, and reliable. For simplicity, the polar diagram shows only a subset of the keywords used for retrieval. Since citation counts fluctuate over time, the retrieval date for related information is set to October 2024.
@@
As illustrated in Fig. 2, we implement a structured process to construct the RiPAMI database, ensuring that the selected articles are review papers relevant to the pattern recognition field. The process begins with searching with keywords derived from the scopes of relevant journals and conferences. These keywords are then formatted into query strings for the arXiv API as follows: (ti:“review” OR ti:“survey”) AND (ti:“keyword.lower()” OR abs:“keyword.lower()”). This structure specifically retrieves articles whose titles or abstracts include both the terms “review” or “survey” and the specified keywords. Additionally, we apply regular expressions to check that the abstracts of returned articles include the relevant keywords, ensuring that only appropriate review articles enter the RiPAMI database. To avoid potential copyright and licensing issues, papers are retrieved and downloaded using the arXiv’s API. Calls to the API are made by means of HTTP requests to a certain URL. The responses will be parsed and stored in an SQL-based database. Through this process, we collect a total of 4,106 samples. To further refine our database, we employ a double-checking phase. This involves a GPT-based filter to preliminarily classify papers, followed by a manual validation step to remove any noisy or irrelevant entries. Lastly, we supplement the remaining 3099 entries with additional metadata, including the publication date, citation counts, and reference details. As mentioned in Sec. 2.3.1, we suggest enriching the meta-data of papers by leveraging a combination of disparate data sources. Considering the potential legal risks of crawling to obtain academic data from Google Scholar, Semantic Scholar API was employed to obtain additional meta-data, such as citation and reference details which are not provided by the arXiv API. The final SQL-based snapshot encompasses a wide range of data, including ID, title, publication date, citations, and more, facilitating efficient data retrieval and statistical analysis.
2.4 Bibliometric Indicators for Reviews
Literature reviews reflect the state of their respective fields, highlighting current research priorities and offering a glimpse into emerging trends. Despite their significance, systematic methods for evaluating literature reviews remain underdeveloped. By introducing quantitative tools, bibliometric methods offer a promising solution, enabling researchers to objectively analyze literature reviews and uncover key patterns and trends. Bibliometrics is a research field focused on the use of quantitative methods and statistical analysis to evaluate various aspects of scholarly publications [15]. These methods and analysis, including widely recognized metrics like citation counts and impact factors, offer researchers valuable insights in their daily scientific work, aiding them in making prompt and informed decisions. However, most existing bibliometric methods suffer from several limitations including unfair comparison, misuse, and manipulation as introduced in the “San Francisco Declaration on Research Assessment” (DORA) [30] and the Leiden Manifesto [41]. Notably, many of these metrics were not specifically designed for evaluating literature reviews, further limiting their applicability in this context.
2.4.1 Topic Normalized Citation Success Index (TNCSI)
To fairly compare the external influence of literature reviews across different fields within an acceptable cost, we propose guiding Large Language Models (LLMs) to generate the topic key phrase and then calculating the success index of the current review within the same topic (see Appendix A.1 for details). The $TNCSI$ estimates the impact of research publications within an LLM-generated topic by normalizing citation counts to a scale between 0 and 1. The formula for calculating the $TNCSI$ is:
[…]
where $\lambda$ is the scale parameter obtained through maximum likelihood estimation. Detailed steps for estimating $\lambda$ and constructing the probability density function are provided in the appendix.
The $TNCSI$ demonstrates favorable mathematical properties and interpretability. First, the $TNCSI$ algorithm employs maximum likelihood estimation to convert the probability mass function into a probability density function. This process ensures that, in theory, the $TNCSI$ differentiates between papers with distinct citation counts, avoiding the assignment of identical values to them. Second, the $TNCSI$ possesses physical significance, representing the probability that a specific paper’s citation count surpasses that of any other paper on the same topic. For example, a paper with a $TNCSI$ of 0.5 means it has more citations than half of the papers within the same topic. Furthermore, the calculation of $TNCSI$ is computationally efficient, as it avoids requiring a complete or precise ranking of all related papers. Instead, it provides a reliable estimation of the probability that the given paper surpasses others in terms of citation count within the same topic.
2.4.2 Impact Evolution Index (IEI)
The $IEI$ is defined as the average slope across $l$ distinct points on a $n=l-1$ degree Bézier curve representing the citation trend over time. The formula for calculating $IEI_{L_{l}}$ is:
$IEI_{L_{l}}=\sum_{a=0}^{l-1}\frac{(y_{i}/x_{i})}{l},$ (2)
where $l$ represents the number of months observed. $x_{i}$ and $y_{i}$ are the components of the tangent vector at the $a$-th point on the Bézier curve, indicating the magnitude along the $x$- and $y$-axes, respectively. More details can be found in Appendix A.2
2.4.3 Reference Quality Measurement (RQM)
$RQM=1-e^{-\beta\cdot e^{-(1-ARQ)\cdot S_{mp}}}.$ (3)
In this equation, $\beta$ represents the shift parameter, $ARQ$ denotes the average reference quality, while $S_{mp}$ refers to the median semester count of the reference age, indicating the time from the publication date of the references to the issuance of the review (which will be detailed in the Appendix A.3).
Beyond simply assessing the average impact of references, the $RQM$ integrates references’ timeliness and its influence on review quality. This metric addresses the tendency among authors to rely heavily on classic works, which may no longer fully align with current research trends. Additionally, the adjustment coefficient in $RQM$ is calibrated through statistical and optimization algorithms, tailored to the unique characteristics of each research discipline. This data-driven approach minimizes reliance on heuristic parameter settings and increases RQM’s adaptability across various academic disciplines.
2.4.4 Review Update Index (RUI)
$RUI=p\cdot CDR+q\cdot RAD.$ (4)
Here, $p$ and $q$ are weighting coefficients, set to 10 and 5, respectively. Detailed definitions and calculations of the $CDR$ and $RAD$ are provided in Appendix A.4.
2.5 Information Extraction for Reviews
Given that this study aims to identify the common features of reviews within the field, solely relying on manual inspection would be both impractical and resource-intensive. To further investigate the features of reviews in the PAMI field, we employed LLM-based information extraction techniques to analyze the review content and parse the response to structured results. The proposed approach converts unstructured text into structured data, facilitating further analysis and interpretation. Details regarding the information extraction method can be found in Appendix B.
3 Classical Review of Reviews in PAMI
This chapter provides a classic, narrative overview of the PAMI field, aiming to identify common features shared by existing reviews and laying the foundation for the subsequent statistical analysis.
3.1 Popular Structure of Reviews in PAMI
The review’s structure is alternatively referred to as the framework of the review. It is the outline that authors consider when starting to conduct the survey. A well-designed structure is believed to enhance the paper’s readability and facilitate the reader’s comprehension of the paper’s concepts and knowledge. Typically, this framework encompasses foundational parts or sections such as the abstract, introduction, methodology, discussion, and conclusion, each fulfilling a specific function.
3.1.1 Abstract
The abstract is the most important part of a paper, as it typically summarizes the core content, helping readers quickly grasp the main points and conclusions of the article. An abstract may vary in its level of detail, but a structured abstract improves clarity by dividing the content into clearly labeled sections, enabling readers to quickly understand the main aspects of the study. A structured abstract consists of labeled sections such as Background, Objective, Methods, Results, Conclusion/Discussion, and possibly Limitations [55, 12]. For example, Sariyanidi et. at. [103] structured their abstract by providing the background on facial affect, outlining the persistent challenges in the field, and summarizing their research methods, results, and discussion.
3.1.2 The Beginning Part
Similar to the research article, the introduction section of the literature review usually includes contextualizing the research topic, identifying the knowledge gap, defining the scope and objectives, and laying the foundation knowledge for wider audiences. “Introduction” usually lies at the very beginning of the paper. Most reviews first introduce the definition of the topic to acquaint readers with a basic understanding of the field. For instance, the term “Named Entity Recognition” may be unfamiliar to many scholars outside the field of natural language processing. A clever introduction might provide the origins of the terminology and inform the readers what is named entity recognition, as has been done in Ref [88]. A brief definition of the field in the introduction is also welcomed for some relatively popular fields, e.g., the literature review [102] examines the concept of image categorization in the first sentence.
Preliminaries and problem formulations are also popular with readers, as they both provide profound background knowledge. Given that the Gumbel-max involves considerable mathematical concepts and calculations, a “Preliminaries” section in the review of the Gumbel-max trick [44] serves the purpose of providing background information and basic understanding related to the Gumbel-max trick.
For systematic reviews, a clear statement regarding the review methodology (e.g., literature search strategy and criteria for inclusion and exclusion) is essential [92, 54]. Such a statement helps to minimize potential bias and enables other researchers to replicate the study’s findings. Memon et. at. [82] dedicate a section to detailing their review process, including subsections on the review protocol, inclusion and exclusion criteria, search strategy, study selection process, quality assessment criteria, and data extraction and synthesis, ensuring transparency and rigor in their methodology.
3.1.3 The Middle Part
The middle part of the survey paper, also known as the review part, presents a detailed examination of the relevant research studies, methodologies, findings, and theories related to the chosen theme of the review. Beyond that, The middle part synthesizes the information from related studies to provide a cohesive and integrated understanding of the topic. This synthesis not only aligns the disparate studies but also evaluates their contributions and the interrelations among them. It often includes comparative analysis, highlighting similarities and differences in approaches, and may identify limitations in the existing research that warrant further investigation.
To understand more comprehensively how the middle part is organized and synthesized, we examine a large number of review papers in the field of PAMI and propose a typology for organizing the review section based on these analysis. By analyzing the writing styles and structures of the review part, we categorize them into three main types: method clustering, challenge-oriented, and others.
Similar to the mapping or scoping review, the method clustering-based review part refers to grouping methods according to their technical characteristics and presenting them in separate sections or subsections. Authors need to identify different sections depending on the topic of the article and arrange the same type of methods as closely as possible. Featuring a clear and well-organized article structure, reviews with method clustering-based parts are widely favored by researchers. The reader can gain a comprehensive understanding of the typology, details, and advancements of technology within a specific field. However, such a type of literature review is less tailored for readers who are completely unfamiliar with the subject. This is because readers who lack basic knowledge of the field may not understand the connection between the methods mentioned in different sections, or are even confused to select a proper algorithm for a certain task. By way of illustration, Minaee et al. conducted a comprehensive survey on deep learning-based image segmentation models [85] in a method clustering-based manner. They grouped the models into 10 categories based on the adopted model architectures, such as fully convolutional models, encoder-decoder-based models, and attention-based models. In this way, the reader can quickly gain an understanding of the typologies of deep learning-based image segmentation models and how they differ from each other.
Known for its practicality, the challenge-oriented review part focuses on presenting research methods and findings that address a specific challenge or task. For this type, each section or subsection focuses on a specific challenge or task and centers the discussion on how to address that challenge. The references cited within each section are primarily directly related to addressing that challenge to provide support and evidence. Challenge-oriented reviews are well-suited to application-oriented fields, where readers at different levels of expertise can easily and quickly seek out appropriate solutions within the literature to address the problem they are facing. Despite the advantages, challenge-oriented reviews may fail to provide a comprehensive elaboration of fundamentals and methods details. Bandini et al. divided their survey about egocentric vision hand analysis into several sections including hand segmentation, hand detection, and hand identification [8]. Each section contains multiple subsections, and each of the subsections investigates a certain challenge. (e.g. subsections entitled “Robustness to Illumination Changes” and “Lack of Pixel-Level Annotations” were arranged in the “Hand Segmentation” section ).
In addition to those two types of reviews, a hybrid middle part strikes a fair balance between comprehensiveness and practicality. Typically, the hybrid middle part refers to a challenge-oriented framework incorporating a method clustering-based sub-framework. One advantage of a hybrid review is that it simultaneously offers multiple solutions to a certain problem along with the classification of the respective methodologies. An example is the study carried out by Guo et al. [37] in which authors divide sections by tasks and introduce similar approaches group by group.
It is worth noting that the primary goal of the typology is not to exhaustively categorize every type of review but to help readers and automated review generation systems better understand the common review structures within the PAMI field. Consequently, certain specialized types of reviews, such as meta-analysis, may fall outside the scope of our typology.
3.1.4 The Ending Part
The subsequent ending part serves as a succinct conclusion to the reviewed studies. Within this section, the authors prefer to sum up the key findings to answer the question in the beginning part. It usually highlights both the advantages and disadvantages of reviewed literature to conclude research gaps and suggest potential avenues for future research.
One of the main objectives of the concluding part is to provide a relatively concise summary of the main insights derived from the reviewed studies and highlight their significance and relevance to the research topic. Some literature reviews conduct comprehensive analysis and synthesis, which in turn run up to a considerable number of pages. The paper [63] consists of 51 double-column pages in the main body, which poses a challenge for readers who may not have the time to peruse it thoroughly. Fortunately, this paper also includes a concise summary which provides readers with an efficient understanding of the content and allows them to navigate to relevant sections of interest.
When discussing further in the ending part, the authors may attempt to identify gaps and point out future directions through the analysis of existing literature. Some of the papers present the gaps and future directions separately. Hassain et al. [45] first discussed major challenges of multi-view video summarization such as lack of synchronization, instability of camera, and crowded scenes individually. Where there are challenges, there are future research directions. In addition to challenges, the authors also provide recommendations and future directions from various perspectives including models, benchmark datasets, and agents-based MVS, etc. Conversely, some papers choose to combine discussions of current issues with emerging research trends, as seen in the ’Future Directions’ section of Ref[135].
3.2 Review of Reviews in Various Sub-fields in PAMI
In this section, we offer a subjective analysis of reviews from diverse fields. Through this narrative approach, we aim to examine the composing characteristics shared by reviews across various fields.
3.2.1 Computer Vision
3.2.2 Natural Language Processing
Sentiment Analysis focuses on determining the sentiment or emotion expressed in text, such as positive, negative, or neutral. Yadav et al. introduce the process of gathering and analyzing people’s opinions and sentiments from various sources such as social media platforms and blogs in their paper [134]. The paper evaluates and compares different approaches used in sentiment analysis, with a focus on supervised machine learning methods like Naive Bayes and SVM algorithms. The common application areas of sentiment analysis and the challenges involved in accurately interpreting sentiments are also reported. The paper by Yue [140] categorizes and compares a large number of techniques and methods from three different perspectives: task-oriented, granularity-oriented, and methodology-oriented. It also explores different types of data and advanced tools for research, highlighting their strengths and limitations.
3.2.3 Others
3.3 Summary
From the discussion above, it is clear that in the field of PAMI, most authors exhibit similar preferences when composing reviews. Common practices include establishing taxonomies, highlighting future challenges and research directions, offering foundational background knowledge, summarizing the performance of various methods on standardized datasets, and examining the real-world applications of algorithms. These preferences will serve as the content features of the review and will be further statistically analyzed in the next section.
4 Statistical Analysis of Reviews in PAMI
In this section, we perform a statistical analysis of reviews in the RiPAMI database, with the goal of uncovering statistical patterns of reviews across various fields of PAMI.
4.1 Meta-data Statistics of Reviews in PAMI

### v6

# Penghai Zhao, Xin Zhang, Jiayue Cao

@@ Abstract
The rapid growth of research in Pattern Analysis and Machine Intelligence (PAMI) has rendered literature reviews essential for consolidating and interpreting knowledge across its many subfields. In this work, we present a comprehensive tertiary analysis of PAMI reviews along three complementary dimensions: (i) identifying structural and statistical regularities in existing surveys; (ii) developing quantitative strategies that help researchers navigate and prioritize within the expanding review corpus; and (iii) critically assessing emerging AI-generated review systems. To support this study, we construct RiPAMI, a large-scale database containing more than 3,000 review articles, and combine narrative synthesis with statistical analysis to capture structural and content-level features. Our analyses reveal distinctive organizational patterns as well as persistent gaps in current review practices. Building on these insights, we propose practical, article-level strategies for indicator-guided navigation that move beyond simple citation counts. Finally, our evaluation of state-of-the-art AI-generated reviews indicates encouraging advances in coherence and organization, yet also highlights enduring weaknesses in reference retrieval, coverage of recent work, and the incorporation of visual elements. Together, these findings provide both a critical appraisal of existing review practices and a forward-looking perspective on how AI-generated reviews can evolve into trustworthy, customizable, and transformative complements to traditional human-authored surveys.
keywords:
AI for research, literature review, umbrella study, AI-generated review, bibliometrics
@@ 1 Introduction
In both natural and artificial systems, entropy—the degree of disorder—tends to increase continuously. A comparable phenomenon occurs in the realm of human knowledge. As research outputs grow rapidly, the knowledge system accumulates scattered and overlapping information, which can lead to redundancy and inefficiency in creating new insights. To counter this trend, literature reviews function as an essential organizing mechanism. Much like gravity that shaped the early universe from dispersed particles into coherent structures, reviews gather fragmented research findings and arrange them into a meaningful whole. A literature review therefore serves not only to demonstrate familiarity with the academic work on a given topic but also to position that work within a broader context. By compiling and synthesizing the most relevant publications, it provides a clear and comprehensive overview of a field and supports the efficient advancement of knowledge.
@@
Figure 1: Annual Publication Trends of Literature Review in the Field of PAMI: A notable rising trend can be observed since 2015, reflecting an increasing scholarly focus and growing recognition of the importance of review articles in synthesizing the state of research in PAMI. The data is collected from the Google Scholar search engine.
@@ 1.1 Research Question and Contribution
To provide a coherent exploration of literature reviews in the PAMI field, this study is organized around three interrelated research questions that progress from current understanding, to practical strategies, and finally to future prospects:
  • RQ1: What structural conventions and statistical patterns characterize literature reviews in the PAMI field?
    To address this question, we construct the RiPAMI database containing more than 3,000 review papers and apply both narrative synthesis and statistical analysis. Specifically, we analyze section structures, authorship patterns, reference and citation distributions, as well as visual element usage. Using LLM-based information extraction, we further quantify six representative content features (e.g., taxonomy, PRISMA, benchmark) and investigate their prevalence across sub-fields and temporal trends. This yields a comprehensive mapping of structural and statistical practices in PAMI reviews.
  • RQ2: Given these observed patterns, how can quantitative bibliometric indicators be designed and applied to help researchers efficiently navigate and select among the growing number of reviews?
    To answer this question, we design four novel, real-time, article-level, and field-normalized bibliometric indicators: the Topic Normalized Citation Success Index (TNCSI), the Impact Evolution Index (IEI), the Reference Quality Measurement (RQM), and the Review Update Index (RUI). We validate these indicators using the RiPAMI database, examining their mathematical properties, interpretability, and correlations with review impact. We further illustrate their practical utility by showing how they support indicator-guided navigation, enabling researchers to efficiently screen and prioritize reviews beyond simple citation counts.
  • RQ3: When applied to emerging AI-generated reviews, what strengths and weaknesses can be observed compared with human-authored surveys, and what do these observations imply for their reliability and practical utility?
    To answer this question, we conduct a systematic evaluation of several state-of-the-art automated review generation systems. Our analysis considers their pipeline design (e.g., intention analysis, retrieval, synthesis, report generation), reference selection strategies, structural organization, and use of visual elements. The results reveal clear patterns: recent systems are able to generate coherent and reasonably organized reviews, sometimes enriched with figures or taxonomies, yet they still suffer from critical shortcomings. These include a tendency to over-rely on highly cited but outdated references, limited capacity to recognize and integrate very recent work, and insufficient incorporation of explanatory visuals or appraisal criteria, etc. Such findings not only highlight the current gap between automated and human-authored reviews, but also underscore the importance of improving reliability, transparency, and customization if AI-generated reviews are to become practically useful in scholarly practice.
@@ 1.2 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 briefly introduces the criteria for literature screening, methods for the RiPAMI construction, and formulas for metric calculations. Section 3 integrates narrative synthesis with statistical examination, providing both a qualitative account of structural and content patterns and a quantitative mapping of meta-data features across existing surveys. Section 4 further discusses how the proposed metrics can assist in efficiently selecting the proper reviews. The characteristics of human-authored versus AI-generated literature reviews are discussed in Sec. 5. In Sec. 6, we explore the challenges and future prospects of literature reviews. Finally, the paper concludes in Sec. 7.
@@ 2.1 Scope
This study examines all fourteen types of review articles as classified in Grant’s comprehensive typology [33], including narrative review, systematic review, state-of-the-art review, etc. While this research is technically a tertiary or ``umbrella'' study, we opted for the more commonly understood term ``literature review'' to improve clarity and acceptance within the PAMI field. In addition, although the quantity of reviews is considerably smaller compared to that of normal papers, it remains impractical to analyze reviews within every field. Therefore, this paper will focus only on reviews in the PAMI field.
2.2 Literature Selection Criteria for Narrative Reviews
To support the analysis in Section 3.1, we establish specific criteria for selecting literature included in the narrative review. The search process draws on Google Scholar, Semantic Scholar, IEEE Xplore, and ScienceDirect, covering publications from major publishers such as IEEE, Elsevier, Springer, and others.
The filtering procedure follows three steps. First, we compile 106 search keywords derived from the scopes of leading journals and conferences in the PAMI domain, including terms such as speech recognition, optical character recognition, and self-supervised learning. Second, we require that a paper’s title contains either “survey” or “review,” and that the chosen keyword appears in its title or abstract. Finally, we exclude preprints not published in journals or conference proceedings, ensuring that all selected articles are peer-reviewed.
These criteria provide a focused and high-quality set of review articles, which serve as the basis for the narrative analysis presented in Section 3.1.
@@ 2.3.1 Data Source
Table 1: Comparisons between Various Sources: ``Counts only'' in the citations column means that the source only records the citation counts, while ``Complete'' signifies that the data source provides a complete list of citations.
@@
Table 1 compares various most commonly used data sources. However, none of these sources is perfect. For example, arXiv is a free distribution service and an open-access archive for millions of scholarly articles in various fields. However, the arXiv data source fails to provide the publication venue, citations, and references. Semantic Scholar seems promising, but it suffers a lower update rate than arXiv and a narrower search scope than Google Scholar. Google Scholar is an online search engine that indexes scholarly literature from a wide range of disciplines. Although Google Scholar uses powerful automated programs to retrieve files for inclusion in its search results, it still faces challenges in many areas. Beel [11, 9, 10] argues that Google Scholar places a high weight on citation counts in its ranking algorithm and has therefore been criticized for exacerbating the Matthew effect [85]. Moreover, the citation counts displayed on Google Scholar are subject to manipulation by complete nonsense articles indexed on Google Scholar (e.g. citations from AI-generated pre-print papers published on arXiv should have been ignored). Therefore, a promising engineering solution is to leverage the strengths of various approaches to overcome the weaknesses of each approach, as will be detailed in the next subsection.
@@ 2.3.2 Database construction
To ensure reproducible experiments and prevent overburdening the server, we construct an SQL-based database dubbed RiPAMI (Reviews in Pattern Analysis and Machine Intelligence, pronounced as /ri:pæmi/). This database stores information related to the paper such as title, abstract, date of publication, venue, citation counts, and reference details, etc.
@@
Figure 2: A Schematic Overview of the Database Construction Process: From initial keyword selection to the final SQL-based RiPAMI snapshot, three key steps are implemented to ensure the data in RiPAMI is clean, accurate, and reliable. For simplicity, the polar diagram shows only a subset of the keywords used for retrieval. Since citation counts fluctuate over time, the retrieval date for related information is set to October 2024.
@@
As illustrated in Fig. 2, we implement a structured process to construct the RiPAMI database, ensuring that the selected articles are review papers relevant to the pattern recognition field. The process begins with searching with keywords derived from the scopes of relevant journals and conferences (check  D for the entire keywords list). These keywords are then formatted into query strings for the arXiv API as follows: (ti:``review'' OR ti:``survey'') AND (ti:``keyword.lower()" OR abs:``keyword.lower()''). This structure specifically retrieves articles whose titles or abstracts include both the terms ``review'' or ``survey'' and the specified keywords. Additionally, we apply regular expressions to check that the abstracts of returned articles include the relevant keywords, ensuring that only appropriate review articles enter the RiPAMI database. To avoid potential copyright and licensing issues, papers are retrieved and downloaded using the arXiv's API. Calls to the API are made by means of HTTP requests to a certain URL. The responses will be parsed and stored in an SQL-based database. Through this process, we collect a total of 4,106 samples. To further refine our database, we employ a double-checking phase. Papers are first preliminarily classified using a GPT-based filter, followed by a manual validation step in which four junior annotators conduct the initial screening and two advanced annotators provide the final confirmation to ensure the removal of false positives. Lastly, we supplement the remaining 3,099 entries with additional metadata, including the publication date, citation counts, and reference details. As mentioned in Sec. 2.3.1, we suggest enriching the meta-data of papers by leveraging a combination of disparate data sources. Considering the potential legal risks of crawling to obtain academic data from Google Scholar, Semantic Scholar API was employed to obtain additional meta-data, such as citation and reference details, which are not provided by the arXiv API. The final SQL-based snapshot encompasses a wide range of data, including ID, title, publication date, citations, and more, facilitating efficient data retrieval and statistical analysis.
2.3.3 Meta-data Statistics of RiPAMI